# Epidemic-induced local awareness behavior inferred from surveys and genetic sequence data

### v2

@@ Introduction
Figure 1: (a) The MASZK Hungarian telephone survey, with 1000 participants in each of the 9 months, shows that the mean local awareness score (in blue) remains relatively constant throughout the recording period, except during the Omicron wave, when the score drops. The government-imposed preventive measures (mask wearing, in yellow, and social distancing, in green) show a different temporal pattern. The daily number of cases (normalized by 170) are shown in red. (b) Our proposed pipelines to generate synthetic (blue) and process real genetic sequence data (grey) to compute collision clusters, superspreading events (SSEs), and finally Event Containment Scores (ECSs) – a proxy measure for local awareness behavior.
@@
While genetic data raises relatively minor privacy concerns [19], and it is unparelleled in terms of availablity at the individual level, inferring behavioral information from genetic sequences is a challenging task. In phylodynamics [20, 21], human behavior is typically estimated based on the phylogenetic tree reconstructed from the observed sequences [22]. However, current tree reconstruction methods have a number of limitations. First, traditional methods are computationally intensive and it is difficult to scale them to datasets with more than a few thousand sequences [23, 24]. Since the COVID-19 pandemic, there has been significant progress in developing more scalable methods [25], and releasing publicly available trees for further analysis [26, 27]. However, processing millions of SARS-CoV-2 genetic sequences remains a challenge [28], and the publicly shared pre-computed trees do not have the same coverage as the Global Initiative on Sharing All Influenza Data (GISAID) dataset, which contains over 16 million SARS-CoV-2 genetic sequences, with a 5-15% sequencing rate in several countries [29]. Second, working with general-purpose methods or highly pre-processed datasets can significantly lower the statistical power of our results, especially since previous methods were not optimized to measure local awareness behavior. Instead, we process this new dataset of unprecedented size by focusing on a simple and tractable statistic that does not require the reconstruction of the phylogenetic tree – the size distribution of the clusters of identical genetic sequences over time. Similar tree-free methods with different applications have been recently proposed by [30, 31, 32]. In essence, we break up the global epidemic into thousands of sub-epidemics with identical genetic code to infer patterns of local awareness. Since each sub-epidemic contains only very noisy information about general local awareness patterns in the population, we focus on one of the most robust features of the dataset: superspreading events.
@@ Results
We detect superspreading events in each collision cluster by tracking unexpectedly large increases in their size after proper normalization (see Methods). Our superspreading event detection method is closely related to previous thresholding approaches [32, 39], requires only minor preprocessing. The detected events agree with our intuition after visual inspection (Figure 2 (b)) and a more in-depth analysis based on location metadata in Supplementary Section A. Thereafter, we assign Event Containment Scores (ECSs) to each superspreading event by comparing the size of the collision clusters after superspreading events and after appropriately selected baseline events during the same time period (see Methods). Finally, to acquire aggregate descriptions of event containment, we compute the median of $\mathrm{ECS}$ values in each country-variant pair $c$, denoted by $\mathrm{ECS}_{c}$; the output of the pipeline in Figure 1 (b). Intuitively, a positive $\mathrm{ECS}_{c}$ means that superspreading events typically led to smaller collision cluster sizes, and therefore fewer secondary infections than the baselines, i.e. the superspreading events were well-contained (Figure 2 (b), red squares). Similarly, a negative $\mathrm{ECS}$ would suggest superspreading events that were not contained as well as the baselines (Figure 2 (b), blue squares).
@@
We set up a synthetic pipeline (Figure 1 (b)) to generate genetic sequence datasets similar to the GISAID EpiCoV dataset, which we can analyze with our superspreading event detection and ECS assignment pipeline. First, we simulate Susceptible-Infected-Recovered (SIR) epidemics on various synthetic and real networks, then we apply the Jukes-Cantor [43] genetic substitution model on the resulting infection tree to produce genetic sequence data (see Methods). To model the combined effect of not all infectious individuals being identified (detection rate), and not all identified individuals being sequenced (sequencing rate), we randomly subsample the generated sequences with probability $p$. Finally, we compute the $\mathrm{ECS}_{c}$ values as before, with $c$ denoting the model parameters instead of the country-variant pair.
@@
Figure 3: Event Containment Scores (ECS) computed on genetic sequence data generated from simulated epidemics on synthetic and real networks as a function of (a) the local, (b) the global awareness function parameter, (c) the infection probability and (d) the subsampling probability. When not stated otherwise, all parameters are set to be their default values $\alpha_{l}=0$, $\alpha_{g}=0$, $\beta_{0}=0.15$ and $p=1$. We observe positive $\mathrm{ECS}_{c}$ values in case of local awareness, and noisy $\mathrm{ECS}_{c}$ values near zero if the subsampling probability is low.
In Figure 3, we plot the dependence of $\mathrm{ECS}_{c}$ on the awareness-strength parameters $\alpha_{l}$ and $\alpha_{g}$ and two potential confounding factors: the basic infection probability $\beta_{0}$, and the subsampling probability $p$. The results indicate that $\mathrm{ECS}_{c}$ primarily depends on the parameter $\alpha_{l}$ (Figure 3 (a)). Importantly, we were only able to generate positive $\mathrm{ECS}_{c}$ values with the local awareness model, apart from the noisy $\mathrm{ECS}_{c}$ values near zero for low subsampling probability in smaller networks. This is a strong indication that that the positive $\mathrm{ECS}_{c}$ values are signs of local awareness behavior.
The observation that only local awareness can produce positive $\mathrm{ECS}_{c}$ values has an intuitive explanation. When a superspreading event occurs, there is usually a common trait between the individuals that become infected at the same time; they all tend belong to the same community as the initial infector. It is also likely, that there exist many additional individuals that belong to the same community, but do not become immediately infected. Indeed, reports of early superspreading events during COVID-19 do not report all individuals becoming infected in the communities at the same time [50, 51], and the same is true in simulations, unless the infection probability inside the community is close to 1. If the structure of the contact network remains unchanged after the superspreading event, then these additional community members become infected in the next timestep (week), which causes the number of sequences in the collision cluster to grow, and therefore produces a negative $\mathrm{ECS}_{c}$ value. Note that there are extreme examples of static networks and epidemic parameters that produce a positive $\mathrm{ECS}_{c}$ value. For instance, in a star network with infection probability close to 1, an epidemic from the center node produces a single superspreading event, and then dies out in the next step, resulting in $\mathrm{ECS}_{c}>0$. However, we conclude that besides a few extreme cases, positive $\mathrm{ECS}_{c}$ values – such as the ones observed in the empirical dataset in Figure 4 – are signs of local awareness behavior.
@@
We compute the $\mathrm{ECS}_{c}$ values for all country-variant pairs with at least 15 detected superspreading events during the Delta or the Omicron BA.1 variants in the GISAID EpiCoV dataset (Figure 4 (a)-(b)), and we analyze how these values are related to behavioral metrics and potential confounding factors. Since we only have 5 datapoints in the Omicron BA.1 wave due to data availability, we also performed the same experiment on all Omicron sequences merged together in Figure 4 (c).
Figure 4: Event Containment Scores (ECS, blue) and containment health index (green) in European countries with at least 15 detected superspreading events in the (a) Delta, (b) Omicron BA.1 variants, and (c) when all Omicron variants are merged. All bar plots show median values and corresponding confidence intervals, and datapoints with a confidence interval larger than 3 are filtered out. Grey background signifies a statistically significant correlation between $\mathrm{ECS}_{c}$ and the containment health index (Table 1).
@@
Spearman-r statistic 0.800 0.800 1.000 -0.867 $0$0.107
Spearman-r p-value $0.010^{*}$ 0.104 $0.000^{*}$ $0.002^{*}$ 1.000 0.819
Table 1: Spearman-r statistics and p-values computed between the $\mathrm{ECS}_{c}$ values and the exogenous variables (containment health index plotted in Figure 4 (a)-(c) and sequencing rate plotted in Supplementary Figure B.3). Significant p-values are marked with a star and with grey.
Figure 4 (a)-(b) shows statistically significantly positive containment scores for Germany in the Delta wave and Germany, Slovenia and Belgium during the Omicron BA.1 wave – a sign of local awareness behavior established in the previous section. To understand the factors that could explain the variability between the observed $\mathrm{ECS}_{c}$ values, we compute the sequencing rate, the attack rate and the Containment Health Index (CHI) in each country-variant pair (see Methods). CHI is a composite epidemic response measure based on thirteen policy indicators maintained by the Oxford Coronavirus Government Response Tracker (OxCGRT) project, similarly to the stringency index [40]. We plot the CHI in Figure 4 (a)-(c), and we compute the Spearman-r statistic between them and the $\mathrm{ECS}_{c}$ values (Table 1). Interestingly, we find a positive correlation between the $\mathrm{ECS}_{c}$ values and the Containment Health Index, which becomes statistically significant in the Delta wave and when we merge Omicron waves into a single dataset, suggesting that government policies may also the impact the local awareness behavior we measure.
While we find no significant correlation between the $\mathrm{ECS}_{c}$ values and the attack rate (Supplementary Figure B.3), we do observe a statistically significant negative correlation with the sequencing rate during the Delta wave (Table 1), which could suggest that $\mathrm{ECS}_{c}$ is an artefact of how the data was collected. However, in the Delta wave, sequencing rate and CHI happened to be highly and negatively correlated, potentially because countries aimed to lift the economic burden of strict containment policies by a higher quality sequencing and monitoring project. In the Omicron BA.1 wave and when all Omicron samples are merged, there is no significant correlation between the $\mathrm{ECS}_{c}$ values and the sequencing rate, suggesting that $\mathrm{ECS}_{c}$ measures a behavioral signal instead of confounding effects.
@@
Having validated containment scores in real and synthetic datasets, we return to our motivating research question; whether drops in local awareness behavior can be observed in the genetic sequence dataset during the Omicron BA.1 wave of the COVID-19 pandemic. One approach to answer this question is to compare the variant-aggregated $\mathrm{ECS}_{c}$ scores from Figure 4 between the Delta and the Omicron BA.1 waves. Figure 5 (a) shows that $\mathrm{ECS}_{c}$ values during the Omicron BA.1 wave were lower compared to the Delta wave in Ireland and the United Kingdom, with other European countries either showing no change between the two waves (Belgium), or an increased $\mathrm{ECS}_{c}$ in the Omicron BA.1 wave (Sweden, Denmark, Germany, Slovenia). As opposed to the spatial analysis in Figure 4, the temporal trends in the $\mathrm{ECS}_{c}$ do not seem to be explained by the Containment and Health Index. Figure 5 (b) shows that while the ranking of the $\mathrm{ECS}_{c}$ values and the CHI are still correlated, the median stringency of the policies became more relaxed only in Sweden and in Belgium, and no change can be observed in the case of Ireland and the UK. However, the purpose of ECS values is to measure the impact of local awareness instead of the policy stringency in the country. As an alternative explanation, we highlight the fact that the Omicron BA.1 wave arrived to the UK and to Ireland a few weeks before its arrival to continental Europe, during the late December instead of early January. The extreme changes in mixing behavior during the holiday season may have contributed to the lower containment scores measured in Figure 5 (b).
@@
Figure 5: (a) $\mathrm{ECS}_{c}$ during the Delta and the Omicron BA.1 variants as computed in Figure 4. Datapoints below the dashed ($x=y$) line hint at drops in local awareness during Omicron BA.1 variant. (b) Containment and Health Index (CHI) during the Delta and the Omicron BA.1 variants as computed in Figure 4. (c) $\mathrm{ECS}_{c}$ values computed on a monthly basis in the UK for the Delta, Omicron BA.1 and BA.2 variants. We observe a drop in $\mathrm{ECS}_{c}$ in December 2021 - January 2022 during the Omicron BA.1 wave.
@@
Up until this point, we focused on the $\mathrm{ECS}_{c}$ values, computed as the median of all ECS values for a country-variant pair. However, in the United Kingdom – where thousands of superspreading events are detected across multiple variants – a higher temporal resolution can be achieved by calculating the median of ECS values on a monthly basis. The resulting signal (Figure 5 (c)), obtained purely based on genetic sequence data, shares a remarkable similarly with the Hungarian survey results in Figure 1 (a). Both curves show a relatively stable signal between October 2021 and July 2022, with a smaller drop during November 2021 and a significant drop at the beginning of the Omicron BA.1 wave. While the two signals were collected using vastly different methods and across different countries – rendering direct comparisons inherently challenging – their alignment opens an array of new questions and research directions in behavioral epidemiology. Moreover, the temporal resolution of the ECS signal in the UK underscores the potential of our approach as a new tool to evaluate the impact of local awareness behavior during a pandemic situation.

### v3

@@ Introduction
Figure 1: (a) The MASZK Hungarian telephone survey, with 1000 participants in each of the 9 months, shows that the mean local awareness score (in blue) remains relatively constant throughout the recording period, except during the Omicron wave, when the score drops. The government-imposed preventive measures (mask wearing, in yellow, and social distancing, in green) show a different temporal pattern. For all survey results, we show the mean response for each month, with confidence intervals calculated under the assumption of normality. The daily number of cases (with a rolling-mean of 7 days, normalized by 170) are shown in red. Source data are available in Supplementary Data 1. (b) Our proposed pipelines to generate synthetic (blue) and process real genetic sequence data (grey) to compute collision clusters, superspreading events (SSEs), and finally Event Containment Scores (ECSs) – a proxy measure for local awareness behavior.
@@
While genetic data raises relatively minor privacy concerns [19], and it is unparelleled in terms of availablity at the individual level, inferring behavioral information from genetic sequences is a challenging task. In phylodynamics [20, 21], human behavior is typically estimated based on the phylogenetic tree reconstructed from the observed sequences [22]. However, current tree reconstruction methods have a number of limitations. First, traditional methods are computationally intensive and it is difficult to scale them to datasets with more than a few thousand sequences [23, 24]. Since the COVID-19 pandemic, there has been significant progress in developing more scalable methods [25], and releasing publicly available trees for further analysis [26, 27]. However, processing millions of SARS-CoV-2 genetic sequences remains a challenge [28], and the publicly shared pre-computed trees do not have the same coverage as the Global Initiative on Sharing All Influenza Data (GISAID) dataset, which contains over 16 million SARS-CoV-2 genetic sequences, with a 5-15% sequencing rate in several countries [29]. Second, working with general-purpose methods or highly pre-processed datasets can significantly lower the statistical power of our results, especially since previous methods were not optimized to measure local awareness behavior. Instead, we process this new dataset of unprecedented size by focusing on a simple and tractable statistic that does not require the reconstruction of the phylogenetic tree – the size distribution of the clusters of identical genetic sequences over time. Similar tree-free methods with different applications have been recently proposed by [30, 31, 32, 33]. In essence, we break up the global epidemic into thousands of sub-epidemics with identical genetic code to infer patterns of local awareness. Since each sub-epidemic contains only very noisy information about general local awareness patterns in the population, we focus on one of the most robust features of the dataset: superspreading events.
@@ Results
We detect superspreading events in each collision cluster by tracking unexpectedly large increases in their size after proper normalization (see Methods). Our superspreading event detection method is closely related to previous thresholding approaches [33, 40], requires only minor preprocessing. The detected events agree with our intuition after visual inspection (Figure 2 (b)) and a more in-depth analysis based on location metadata in Supplementary Section A. Thereafter, we assign Event Containment Scores (ECSs) to each superspreading event by comparing the size of the collision clusters after superspreading events and after appropriately selected baseline events during the same time period (see Methods). Finally, to acquire aggregate descriptions of event containment, we compute the median of $\mathrm{ECS}$ values in each country-variant pair $c$, denoted by MECS; the output of the pipeline in Figure 1 (b). Intuitively, a positive MECS means that superspreading events typically led to smaller collision cluster sizes, and therefore fewer secondary infections than the baselines, i.e. the superspreading events were well-contained (Figure 2 (b), red squares). Similarly, a negative $\mathrm{ECS}$ would suggest superspreading events that were not contained as well as the baselines (Figure 2 (b), blue squares).
@@
We set up a synthetic pipeline (Figure 1 (b)) to generate genetic sequence datasets similar to the GISAID EpiCoV dataset, which we can analyze with our superspreading event detection and ECS assignment pipeline. First, we simulate Susceptible-Infected-Recovered (SIR) epidemics on various synthetic and real networks, then we apply the Jukes-Cantor [44] genetic substitution model on the resulting infection tree to produce genetic sequence data (see Methods). To model the combined effect of not all infectious individuals being identified (detection rate), and not all identified individuals being sequenced (sequencing rate), we randomly subsample the generated sequences with probability $p$. Finally, we compute the MECS values as before, with $c$ denoting the model parameters instead of the country-variant pair.
@@
Figure 3: Event Containment Scores and their median values (MECS) computed on genetic sequence data generated from simulated epidemics on synthetic and real networks as a function of (a) the local, (b) the global awareness function parameter, (c) the infection probability and (d) the subsampling probability. For each set of parameters, we simulated $n=200$ independent epidemic processes with different random seeds. Colored intervals show the 25th and 75th percentiles of the ECS values, while black intervals indicate confidence intervals for the median, computed using a normal approximation. Source data are available in Supplementary Data 2. When not stated otherwise, all parameters are set to be their default values $\alpha_{l}=0$, $\alpha_{g}=0$, $\beta_{0}=0.15$ and $p=1$. We observe positive MECS values in case of local awareness, and noisy MECS values near zero if the subsampling probability is low.
In Figure 3, we plot the dependence of MECS on the awareness-strength parameters $\alpha_{l}$ and $\alpha_{g}$ and two potential confounding factors: the basic infection probability $\beta_{0}$, and the subsampling probability $p$. The results indicate that MECS primarily depends on the parameter $\alpha_{l}$ (Figure 3 (a)). Importantly, we were only able to generate positive MECS values with the local awareness model, apart from the noisy MECS values near zero for low subsampling probability in smaller networks. This is a strong indication that that the positive MECS values are signs of local awareness behavior.
The observation that only local awareness can produce positive MECS values has an intuitive explanation. When a superspreading event occurs, there is usually a common trait between the individuals that become infected at the same time; they all tend belong to the same community as the initial infector. It is also likely, that there exist many additional individuals that belong to the same community, but do not become immediately infected. Indeed, reports of early superspreading events during COVID-19 do not report all individuals becoming infected in the communities at the same time [51, 52], and the same is true in simulations, unless the infection probability inside the community is close to 1. If the structure of the contact network remains unchanged after the superspreading event, then these additional community members become infected in the next timestep (week), which causes the number of sequences in the collision cluster to grow, and therefore produces a negative MECS value. Note that there are extreme examples of static networks and epidemic parameters that produce a positive MECS value. For instance, in a star network with infection probability close to 1, an epidemic from the center node produces a single superspreading event, and then dies out in the next step, resulting in $MECS>0$. However, we conclude that besides a few extreme cases, positive MECS values – such as the ones observed in the empirical dataset in Figure 4 – are signs of local awareness behavior.
@@
We compute the MECS values for all country-variant pairs with at least 15 detected superspreading events during the Delta or the Omicron BA.1 variants in the GISAID EpiCoV dataset (Figure 4 (a)-(b)), and we analyze how these values are related to behavioral metrics and potential confounding factors. Since we only have 5 datapoints in the Omicron BA.1 wave due to data availability, we also performed the same experiment on all Omicron sequences merged together in Figure 4 (c).
Figure 4: Event Containment Scores (ECS, blue) and Containment Health Index (CHI, green) in European countries with at least 15 detected superspreading events in the (a) Delta, (b) Omicron BA.1 variants, and (c) when all Omicron variants are merged. Bar plots and black dots mark median values. The number of ECS values corresponding to each Median ECS (MECS) value is shown in Supplementary Table B.1. Colored intervals show the 25th and 75th percentiles of the distribution, while black intervals indicate confidence intervals for the median, computed using a normal approximation. Country-variant pairs with a confidence interval larger than 3 around the MECS values are filtered out. Grey background signifies a statistically significant correlation between MECS and the median CHI values (Table 1).
@@
Spearman’s $\rho$ 0.800 0.800 1.000 -0.867 $0$0.107
$p$-value $0.010^{*}$ 0.104 $0.000^{*}$ $0.002^{*}$ 1.000 0.819
Table 1: Spearman rank correlation coefficients (Spearman’s $\rho$) and corresponding two-sided $p$-values were computed between MECS values and the exogenous variables (Containment Health Index plotted in Figure 4 (a)-(c) and sequencing rate plotted in Supplementary Figure B.3). No correction was applied for multiple comparisons. Significant $p$-values ($p<0.05$) are indicated with a star and shaded in grey.
Figure 4 (a)-(b) shows statistically significantly positive containment scores for Germany in the Delta wave and Germany, Slovenia and Belgium during the Omicron BA.1 wave – a sign of local awareness behavior established in the previous section. To understand the factors that could explain the variability between the observed MECS values, we compute the sequencing rate, the attack rate and the Containment Health Index (CHI) in each country-variant pair (see Methods). CHI is a composite epidemic response measure based on thirteen policy indicators maintained by the Oxford Coronavirus Government Response Tracker (OxCGRT) project, similarly to the stringency index [41]. We plot the CHI in Figure 4 (a)-(c), and we compute the Spearman-r statistic between them and the MECS values (Table 1). Interestingly, we find a positive correlation between the MECS values and the Containment Health Index, which becomes statistically significant in the Delta wave and when we merge Omicron waves into a single dataset, suggesting that government policies may also the impact the local awareness behavior we measure.
While we find no significant correlation between the MECS values and the attack rate (Supplementary Figure B.3), we do observe a statistically significant negative correlation with the sequencing rate during the Delta wave (Table 1), which could suggest that MECS is an artefact of how the data was collected. However, in the Delta wave, sequencing rate and CHI happened to be highly and negatively correlated, potentially because countries aimed to lift the economic burden of strict containment policies by a higher quality sequencing and monitoring project. In the Omicron BA.1 wave and when all Omicron samples are merged, there is no significant correlation between the MECS values and the sequencing rate, suggesting that MECS measures a behavioral signal instead of confounding effects.
@@
Having validated containment scores in real and synthetic datasets, we return to our motivating research question; whether drops in local awareness behavior can be observed in the genetic sequence dataset during the Omicron BA.1 wave of the COVID-19 pandemic. One approach to answer this question is to compare the variant-aggregated MECS scores from Figure 4 between the Delta and the Omicron BA.1 waves. Figure 5 (a) shows that MECS values during the Omicron BA.1 wave were lower compared to the Delta wave in Ireland and the United Kingdom, with other European countries either showing no change between the two waves (Belgium), or an increased MECS in the Omicron BA.1 wave (Sweden, Denmark, Germany, Slovenia). As opposed to the spatial analysis in Figure 4, the temporal trends in the MECS do not seem to be explained by the Containment Health Index. Figure 5 (b) shows that while the ranking of the MECS values and the CHI are still correlated, the median stringency of the policies became more relaxed only in Sweden and in Belgium, and no change can be observed in the case of Ireland and the UK. However, the purpose of ECS values is to measure the impact of local awareness instead of the policy stringency in the country. As an alternative explanation, we highlight the fact that the Omicron BA.1 wave arrived to the UK and to Ireland a few weeks before its arrival to continental Europe, during the late December instead of early January. The extreme changes in mixing behavior during the holiday season may have contributed to the lower containment scores measured in Figure 5 (b).
@@
Figure 5: (a) Median Event Containment Scores (MECS) during the Delta and the Omicron BA.1 variants as computed in Figure 4. Datapoints below the dashed ($x=y$) line hint at drops in local awareness during Omicron BA.1 variant. (b) Containment Health Index (CHI) during the Delta and the Omicron BA.1 variants as computed in Figure 4. (c) MECS values computed biweekly with a 4-week sliding window in the UK for the Delta, Omicron BA.1 and BA.2 variants. Confidence intervals were computed using a normal approximation, and datapoints with a confidence interval larger than 2 are filtered out. We observe a drop in MECS in December 2021 - January 2022 during the Omicron BA.1 wave.
@@
Up until this point, we focused on the MECS values, computed as the median of all ECS values for a country-variant pair. However, in the United Kingdom – where thousands of superspreading events are detected across multiple variants – a higher temporal resolution can be achieved by calculating the median of ECS values biweekly with a 4-week sliding window. The resulting signal (Figure 5 (c)), obtained purely based on genetic sequence data, shares a remarkable similarly with the Hungarian survey results in Figure 1 (a). Both curves show a relatively stable signal between October 2021 and July 2022, with a smaller drop during November 2021 and a significant drop at the beginning of the Omicron BA.1 wave.
Notably, the decline in ECS values coincides with the transition from the Delta to the Omicron BA.1 wave, raising concerns that this trend may reflect the increased transmissibility of the BA.1 variant compared to Delta. However, simulations (Figure 3 (c)) suggest that transmissibility alone has only a minimal impact on ECS values. Furthermore, the BA.2 variant, which also had increased transmissibility relative to BA.1, does not exhibit a similar discontinuity in the ECS signal. A possible alternative explanation is that a temporal drop in ECS could also be driven by spatial variations in behavior. Although our analysis treats the UK as one homogeneous population, it has been reported that the introduction of the BA.1 variant into the UK was initially localized in the London area [53], and the drop in ECS could be a result of a limited ability to engage in awareness behavior in this region. In contrast, the Hungarian dataset was a representative survey, and with the Omicron wave arriving later in Hungary, the introduction of the disease was likely more uniform than in the UK.
While the uncertainties and the differences in the data collection render the direct comparison of the British ECS signal and the Hungarian survey inherently challenging, their alignment opens an array of new questions and research directions in behavioral epidemiology. Moreover, the temporal resolution of the ECS signal in the UK underscores the potential of our approach as a new tool to evaluate the impact of local awareness behavior during a pandemic situation.